question 93435 is similar but i don't understand the process.
question: x^2 + 4x - 12 = 0
no text available and no use of a calculator is allowed.

Answer by checkley75(3666)

X^2+4X-12=0
(X+6)(X-2)=0
X+6=0
X=-6 ANSWER.
X-2=0
X=2 ANSWER.